I need help with this question, my mind is blank. 5(4z)-1/2z =234
Answer by ewatrrr(24785)

5(4z)-1/2z =234
20z - (1/2)z = 234
(40/2)z- (1/2)z = 234
(39/2)z = 234
z = (2/39)234 = 12